OpenStudy (anonymous):

Which is the greatest volume? Choices: 1 liter 1 milliliter 1 deciliter 1 dekaliter

OpenStudy (hoblos):

dekaliter

OpenStudy (2yoututu11):

1 dekalitre = 10 litres = 100 decilitres = 10,000 millilitres 1 litre = 1/10 dekalitre = 10 decilitres = 1,000 millilitres 1 decilitre = 1/100 dekalitre = 1/10 litre = 100 millilitres 1 millilitre = 1/10,000 dekalitre = 1/1,000 litre = 1/100 decilitre

OpenStudy (ash2326):

1 liter=1000 ml 1 ml= 1 ml 1 deciliter=100 ml 1 decaliter= 10000ml

OpenStudy (ash2326):

@JackJones now you tell which is the largest?

OpenStudy (apoorvk):

1 deca=10liters 1deci=0.1liter 1mili=0.001 liter so DEKA it is!!!

OpenStudy (anonymous):

decaliter